# FARM: Functional Group-Aware Representations for Small Molecules

### v2

@@ D.2 Training FG Knowledge Graph Embedding Model for Molecular Structure Representation
In the ComplEx model (Trouillon et al., 2016), each element in a triple $(\mathbf{h},\mathbf{r},\mathbf{t})$ — where $\mathbf{h}$ is the head entity, $\mathbf{r}$ is the relation, and $\mathbf{t}$ is the tail entity — is represented as a complex vector:
@@ D.4 Contrastive Learning: Align SMILES and Structure Representation
$\mathcal{L}_{\text{CL}}=\frac{1}{N}\sum_{i=1}^{N}\max\left(0,\gamma-\text{cosine\_similarity}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})+\text{cosine\_similarity}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})\right)$
@@ D.5 Downstream task finetuning
MoleculeNet tasks are treated as downstream tasks for our FARM model. We freeze all layers of FARM and pair it with a GRU head for both classification and regression tasks. For classification, we use cross-entropy as the loss function, while for regression, we employ mean squared error. The Adam optimizer is applied with a learning rate of $41e-$ 4 and a cosine annealing learning rate schedule with a period of 20 epochs. The training process spans 100 epochs with a batch size of 16, using an 80-10-10 train-validation-test split with scaffold splitting. To address imbalanced datasets, we implement a weighted loss function, assigning a weight of 5 to classes with fewer samples. For each task, we conduct three runs with different train-validation-test splits and report the average and standard deviation of the results.
Appendix E Ablation Study
@@
To assess the effectiveness of each component in our architecture, we conducted a comprehensive ablation study across several MoleculeNet benchmark tasks. The first model, FM_KGE + GAT, utilizes FG knowledge graph embeddings as input for a Graph Attention Network (Veličković et al., 2017) (GAT) to predict molecular properties. Although its performance on these tasks is not the strongest, the model still demonstrates its capacity to learn underlying chemical rules (syntax and semantics) from the data to a certain degree.
@@
The second model, AttentiveFP (Xiong et al., 2019), performs a masked atom prediction task on the molecular graph, predicting atom types such as carbon, hydrogen, oxygen, and nitrogen. Its variation, FG AttentiveFP, shares the same architecture as AttentiveFP, but it predicts both the atom type and the associated functional group. Experimental results indicate that incorporating functional group information significantly improves the model’s performance on downstream tasks.
@@
Table 8 presents the detailed results of the aforementioned models across various MoleculeNet tasks, illustrating the performance of each architecture.
Table 8: Performance of various models across six MoleculeNet tasks. The data is split using a random split into training, validation, and test sets with an 8:1:1 ratio.

### v3

@@ D.2 Training FG Knowledge Graph Embedding Model for Molecular Structure Representation
In the ComplEx model [Trouillon et al., 2016], each element in a triple $(\mathbf{h},\mathbf{r},\mathbf{t})$ — where $\mathbf{h}$ is the head entity, $\mathbf{r}$ is the relation, and $\mathbf{t}$ is the tail entity — is represented as a complex vector:
@@ D.4 Contrastive Learning: Align SMILES and Structure Representation
$\mathcal{L}_{\text{CL}}=\frac{1}{N}\sum_{i=1}^{N}\max\left(0,\gamma-\text{cos}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{pos}})+\text{cos}(\mathbf{h}_{\text{MLM}},\mathbf{h}_{\text{neg}})\right)$
@@ D.5 Downstream task finetuning
MoleculeNet tasks are treated as downstream tasks for our FARM model. We freeze all layers of FARM and pair it with a GRU head for both classification and regression tasks. For classification, we use cross-entropy as the loss function, while for regression, we employ mean squared error. The Adam optimizer is applied with a learning rate of $1e-4$ and a cosine annealing learning rate schedule with a period of 20 epochs. The training process spans 100 epochs with a batch size of 16, using an 80-10-10 train-validation-test split with scaffold splitting. To address imbalanced datasets, we implement a weighted loss function, assigning a weight of 5 to classes with fewer samples. For each task, we conduct three runs with different train-validation-test splits and report the average and standard deviation of the results.
Appendix E Evaluation of FARM on Quantum Mechanical Property Prediction
Table 9: Performance of FARM on quantum mechanical property prediction using the Quantum Mechanical dataset. $R^{2}$ scores are reported with standard deviation over 10 runs.
Class $\mathbf{R^{2}}$ Class $\mathbf{R^{2}}$
HOMOm1 (eV) 0.927±0.01480.927$\pm$ 0.01480.927 ± 0.0148 T5 0.905±0.03150.905$\pm$ 0.03150.905 ± 0.0315
HOMO (eV) 0.895±0.02980.895$\pm$ 0.02980.895 ± 0.0298 S1 0.936±0.01610.936$\pm$ 0.01610.936 ± 0.0161
LUMO (eV) 0.928±0.009480.928$\pm$ 0.009480.928 ± 0.00948 S2 0.948±0.01560.948$\pm$ 0.01560.948 ± 0.0156
LUMOp1 (eV) 0.922±0.02800.922$\pm$ 0.02800.922 ± 0.0280 S3 0.911±0.008490.911$\pm$ 0.008490.911 ± 0.00849
DipoleMoment (Debye) 0.755±0.05320.755$\pm$ 0.05320.755 ± 0.0532 S4 0.919±0.007920.919$\pm$ 0.007920.919 ± 0.00792
LogP 0.995±0.005660.995$\pm$ 0.005660.995 ± 0.00566 S5 0.924±0.01430.924$\pm$ 0.01430.924 ± 0.0143
PrimeExcite (eV) 0.877±0.04910.877$\pm$ 0.04910.877 ± 0.0491 O1 0.839±0.03400.839$\pm$ 0.03400.839 ± 0.0340
PrimeExcite (osc) 0.837±0.02430.837$\pm$ 0.02430.837 ± 0.0243 TDOS2.0 0.868±0.09900.868$\pm$ 0.09900.868 ± 0.0990
T1 0.929±0.005870.929$\pm$ 0.005870.929 ± 0.00587 TDOS2.5 0.877±0.004100.877$\pm$ 0.004100.877 ± 0.00410
T2 0.925±0.01820.925$\pm$ 0.01820.925 ± 0.0182 TDOS3.0 0.815±0.04470.815$\pm$ 0.04470.815 ± 0.0447
T3 0.900±0.01010.900$\pm$ $0$ 0.0101 TDOS3.5 0.895±0.03270.895$\pm$ 0.03270.895 ± 0.0327
T4 0.921±0.01960.921$\pm$ 0.01960.921 ± 0.0196 TDOS4.0 0.919±0.03130.919$\pm$ 0.03130.919 ± 0.0313
TDOS4.5 0.938±0.01830.938$\pm$ 0.01830.938 ± 0.0183
Table 9 presents the results of FARM on the quantum mechanical property prediction tasks from the dataset introduced in [Angello et al., 2024]. FARM demonstrates consistently high performance across a wide range of properties, including difficult-to-predict targets such as Dipole Moment, excited-state energies (S1–S5 and T1–T5), and total density of states (TDOS). These results underscore FARM’s ability to accurately model complex quantum mechanical behaviors by leveraging functional group information. Its success in capturing nuanced molecular interactions further illustrates FARM’s applicability to real-world challenges in chemical property prediction, where understanding the role of functional modules is essential for robust and generalizable predictions.
Appendix F Ablation Study
F.1 FARM Component Analysis
@@
To assess the effectiveness of each component in our architecture, we conducted a comprehensive ablation study across several MoleculeNet benchmark tasks. The first model, FM_KGE + GAT, utilizes FG knowledge graph embeddings as input for a Graph Attention Network [Veličković et al., 2017] (GAT) to predict molecular properties. Although its performance on these tasks is not the strongest, the model still demonstrates its capacity to learn underlying chemical rules (syntax and semantics) from the data to a certain degree.
@@
The second model, AttentiveFP [Xiong et al., 2019], performs a masked atom prediction task on the molecular graph, predicting atom types such as carbon, hydrogen, oxygen, and nitrogen. Its variation, FG AttentiveFP, shares the same architecture as AttentiveFP, but it predicts both the atom type and the associated functional group. Experimental results indicate that incorporating functional group information significantly improves the model’s performance on downstream tasks.
@@
Table 10 presents the detailed results of the aforementioned models across various MoleculeNet tasks, illustrating the performance of each architecture. For these experiments, we used random splitting to divide the downstream datasets into training, validation, and test sets in an 8:1:1 ratio. While random splitting is used consistently across models in this ablation study, scaffold splitting is applied for benchmarking to ensure a fair comparison with other methods.
Table 10: Performance of various models across 6 MoleculeNet tasks. The data is split using a random split into training, validation, and test sets with an 8:1:1 ratio.
@@
F.2 Fragmentation Methods
In this section, we compare the performance of our FG-based fragmentation method with BRICS fragmentation. When using BRICS as the fragmentation method and enriching SMILES with functional group information derived from BRICS (treating each BRICS fragment as a functional group), the resulting vocabulary size is 575,612 (with a minimum frequency of 5), compare to 14,741 with our FG-based fragmentation. This significant difference arises because BRICS tends to generate larger fragments, causing the vocabulary to grow rapidly. This larger vocabulary can make the model harder to train and converge, as it includes a high number of rare tokens. Consequently, the likelihood of encountering out-of-vocabulary tokens during validation and testing increases, which negatively impacts model performance. As shown in Table 4, under the same experimental setup, our FG-based fragmentation method outperforms BRICS fragmentation across all six MoleculeNet tasks.
Appendix G Computational Resources
For the BERT model, even with an expanded vocabulary of approximately 14,000 tokens, the model size remains smaller than the English BERT vocabulary and does not require additional computational resources compared to training a standard BERT model. Each BERT model, with around 100 million parameters, was trained for 20 epochs on a machine with an NVIDIA A100 GPU (40GB memory), 256GB of RAM, and a 32-core AMD EPYC 7502 CPU. Training each model took approximately 6 hours. The experiments were conducted on an internal computing cluster, and no additional compute was required beyond the primary experiments included in the paper.
For other models, such as those used for ComplEx and link prediction GNN, the computational requirements are relatively lightweight, and the same machine used for BERT training was sufficient without the need for additional resources.